### v1

@@ 4.5 Efficiency
Table 4 shows our time and memory requirements for training and inference compared to the baselines. Compared to pixelSplat, we have a slightly faster encoder and only a 1ms slower rendering stemming to 68% from the convolutional decoder and the remaining 32% from splatting additional feature channels. Therefore, we maintain the real-time rendering capabilities of 3D Gaussian splatting despite introducing a generative model. Furthermore, we are memory efficient during both training and inference. Compared to GeNVS, we are orders of magnitude faster in inference and training.
@@ 4.6 Ablations
We conduct an ablation study for the extrapolation setting on CO3D hydrants with the results given in Table 3. Interestingly, while performing much worse in FID and KID with a deterministic version omitting the variational formulation as well as the GAN loss, we obtain state-of-the-art results for PSNR and SSIM. This shows that there is a trade-off between conservative and therefore blurry reconstructions favored by the classical metrics and realistic and detailed novel views rewarded by FID and KID. Without latent and/or RGB skip connections, we perform slightly worse in most metrics validating our design choice of leveraging the rendered features and colors at multiple resolutions. Overall, we find the best balance by using our generative approach together with latent and RGB skip connections. We attribute that to the fact that the decoder has global and local context when interpreting features for image synthesis, which helps generating consistent images.
@@
Table 4: Efficiency comparison. Our method renders roughly 2000 times faster and is much less expensive to train than the state-of-the-art generative model for novel view synthesis GeNVS [8]. Compared to the state-of-the-art regression-based approach, latentSplat entails a negligible increase in resource requirements in turn for significantly better image quality.
@@ 5 Conclusion
We presented latentSplat, a method that successfully combines the strengths of regression-based approaches with the power of a lightweight generative model to handle uncertainty. Our approach achieves state-of-the-art image quality in novel view synthesis while providing the highest perceptual similarity to the ground truth. Compared to previous generative approaches, latentSplat is much faster and more scalable, enabling real-time rendering in large resolutions.
@@
Appendix
In this supplementary material, we give details regarding architecture and evaluation in Sections A and B. Section C deals with the evaluation of 3D consistency using videos. We present visualizations of intermediate results as well as model uncertainty underlining the advantages of latentSplat as a fast generative approach in Section D. Finally, we provide additional qualitative results in Section E. Please also refer to the project website222geometric-rl.mpi-inf.mpg.de/latentsplat/ for results in motion.
@@ A Architecture Details
Loss Our loss can be split into auxiliary losses, reconstruction losses, and a generative loss. We start the training with auxiliary losses only for the first 100k iterations to avoid fitting features to Gaussians in wrong positions. Furthermore, we delay the generative loss until 125k iterations to prevent an overly strong discriminator. In order to account for scale and relevance, the individual terms have different weights. As reconstruction losses, the weights of L1 and LPIPS are $\lambda_{1}=1$ and $\lambda_{2}=1$, respectively. The mean squared error and LPIPS for the axiliary rendering are weighted by $\lambda_{3}=10$ and $\lambda_{4}=0.5$. Finally, we adopt the adaptive weighting of the generative loss from LDM [40] by leveraging the ratio of gradient norms from reconstruction and generative losses to the last layer of the decoder.
@@ B Evaluation Details
C 3D Consistency
@@ D Additional Visualizations
E Additional Results
We provide additional qualitative results for interpolation and extrapolation on CO3D [37] and RealEstate10k [56] in Fig. 8 to 13, and in the supplemental video.

### v2

@@ 4.6 Efficiency
Table 5 shows our time and memory requirements for training and inference compared to the baselines. Compared to pixelSplat [10], our encoding is slightly faster while the rendering is only a 1ms slower stemming to 68% from the convolutional decoder and 32% from splatting additional feature channels. Hence, we maintain the real-time rendering of 3D Gaussian splatting despite introducing a generative model. Furthermore, we are memory efficient during both training and inference. Compared to GeNVS [9], we are orders of magnitude faster.
@@ 4.7 Ablations
We conduct an ablation study for view extrapolation on CO3D hydrants with the results given in Table 4. Interestingly, while performing much worse in FID and KID with a deterministic version omitting the variational formulation as well as the GAN loss, we obtain state-of-the-art results for PSNR and SSIM. This shows the trade-off between conservative and therefore blurry reconstructions favored by classical metrics and realistic and detailed novel views rewarded by FID and KID. Overall, we find the best balance in metric scores by using our generative approach together with multi-resolution latent and RGB skip connections. We attribute that to the decoder having global and local context when interpreting features for image synthesis, which helps generating consistent results.
@@
Table 5: Efficiency comparison. Our method renders roughly 2000 times faster and is much less expensive to train than the state-of-the-art generative model GeNVS [9]. Compared to the regression-based pixelSplat [10], latentSplat entails a negligible increase in resource requirements in turn for significantly better image quality.
@@ 5 Conclusion
We presented latentSplat, a method that successfully combines the strengths of regression-based approaches with the power of a lightweight generative model to handle uncertainty. Our approach achieves state-of-the-art quality in novel view synthesis from two input images while providing the highest perceptual similarity to the ground truth. Predicted views are 3D consistent, enabling downstream mesh reconstruction. Compared to previous generative approaches, latentSplat is much faster and more scalable, allowing real-time rendering in large resolutions.
Limitations and future work. Since the location of Gaussians is limited to the camera frustums of input views, the 2D decoder inpaints out-of-view areas for extrapolation resulting in 3D inconsistencies. Another limitation is the independent sampling of depth along rays and local appearance from variational Gaussians. Finding a trade-off between capturing the full conditional distribution of reconstructions and sampling efficiency renders potential future work.
@@ A Architecture Details
Loss Our loss can be split into auxiliary losses, reconstruction losses, and a generative loss. We start the training with auxiliary losses only for the first 100k iterations to avoid fitting features to Gaussians in wrong positions. Furthermore, we delay the generative loss until 125k iterations to prevent an overly strong discriminator. To account for different scale, the individual terms have different weights. As reconstruction losses, the weights of L1 and LPIPS are $\lambda_{1}=1$ and $\lambda_{2}=1$, respectively. The mean squared error and LPIPS for the axiliary rendering are weighted by $\lambda_{3}=10$ and $\lambda_{4}=0.5$. Finally, we adopt the adaptive weighting of the generative loss from LDM [50] by leveraging the ratio of gradient norms from reconstruction and generative losses to the last layer of the decoder.
@@ B Evaluation Details
C Comparison with GeNVS
We provide an additional qualitative comparison with GeNVS [9] in Fig. 6. Their code is not available and we are not able to reproduce their results for a fair comparison. We still aim to do a comparison using figures from their paper. Note that there are major differences between the setups of GeNVS and latentSplat such as the number of input views and the version of the dataset (CO3Dv1 vs CO3Dv2). We match their setting as close as possible by creating a test split of all their evaluation examples, finding closest poses, and using the same input view together with a close second view. Although the comparison of single- and two-view 3D reconstruction is not fair, note that GeNVS uses scale normalization and depth cues (c.f. supplementary material C.3) to resolve the scale ambiguity from a single input view. The qualitative comparison shows similar image quality. However, our approach is much faster and efficient during both inference and training, as detailed in Sec. 4.6 of the main paper.
Figure 6: Comparison with GeNVS [9] We qualitatively compare against GeNVS. Note that the setups of both methods differ strongly and that GeNVS code is not available for reproducing results. Thus, this comparison is not fair. We selected the same CO3D test examples and compared against renderings shown in their paper. Both methods generate similar quality while ours is much faster (c.f. Sec. 4.6 main paper).
D 3D Consistency
@@ E Additional Visualizations
F Additional Results
We provide additional qualitative results for interpolation and extrapolation on CO3D [47] and RealEstate10k [80] in Fig. 9 to 14, and in the supplemental video.